OpenStudy (anonymous):

Find all solutions to the equation. 7 sin^2(x) - 14 sin x + 2 = -5

OpenStudy (anonymous):

Please someone help!

OpenStudy (anonymous):

i cant put a screen shot but it looks like this

OpenStudy (anonymous):

@ Mathstudent123

OpenStudy (dumbcow):

solve it like a quadratic

OpenStudy (ranga):

This is a quadratic equation in sin(x). If that is confusing just temporarily make the substitution sin(x) = t Solve the quadratic equation in t. Then put back t = sin(x) and solve for x.

OpenStudy (dumbcow):

factor out a 7 let t = sin x --> t^2 -2t +1 = 0

OpenStudy (anonymous):

so 7+14(t)(x)+2=-5

OpenStudy (anonymous):

or 7T^2(x)-14T(x)+2=-5?

OpenStudy (dumbcow):

the 2nd one...but move the -5 over and factor out a 7